# NoTeeline: Supporting Real-Time Notetaking from Keypoints with Large Language Models

### v1

@@ 5.1. Design Goals
Based on our needfinding study, iterative prototyping and evaluation process, we propose three design goals for NoTeeline.
D1. Micronotetaking as the way to take notes from videos. The interface should enable the swift creation of micronote. Even if the micronote includes grammatical errors and/or is extremely short, the system should be able to comprehend it. While transforming micronotes, it should be aware of the context, i.e: the topic being discussed in the video. The system should enable iterative editing of micronotes. By focusing on micronotes, we want to tackle the difficulty in writing full notes users mentioned (C1).
D2. Maintain consistency with user’s style and preferences. The system should take into account the specific writing styles each user follows and remain consistent with their personal style. At the same time, the final note should exclude information users are not interested in. It should not include any hallucinated information, as desired by the participants in our findings (C2).
D3. Each component is linked to a specific note-taking procedure and shown accordingly. The system should link each component to a specific stage in the note-taking process—creation, editing, and revision. It should allow users to engage selectively with these features based on the current stage, providing ease of use throughout the note-taking journey (Pea, 1987; Nielsen, 2006) (C3).
@@ 5.2. Interface Components
We will now describe the components of NoTeeline built to support our design goals.
@@
We design an onboarding session to get reference examples of the user’s specific style (D2). Users watch three short clips (20-30 seconds each); and provide a keypoint as well as the corresponding full note for each clip (Figure 2). Once completed, the system will store these examples to use for note expansion afterwards. Users need to complete this session only once when they start using NoTeeline for the first time. This session also give users a first-hand experience of what a keypoint would look like and what kind of full note to expect from the system.
5.2.2. Notetaking Layout.
To construct the layout, we took inspiration from the Cornell method (Pauk and Owens, 2010), a popular technique for taking note (Saran et al., 2022; Piolat, 2007). We split NoTeeline layout into three regions – 1) Notes: the region for writing and organizing note; 2) Cues: questions or specific points that are helpful to review the note; and 3) Summary: main takeaway of the note in 5 sentences.
We also incorporated the video viewport on the top of the layout so that users can take notes while watching the video without needing to move between different desktop tabs (C3). It also progressively discloses the features as per users’ current need for note writing, editing or analyzing (D3). Figure 2 illustrates the notetaking layout and its different components.
Figure 3. Note expansion (some texts have been cropped for clarity).
@@
Micro Note Expansion. To construct the full note, we have an atomic expansion operation. For expansion, we utilize these pieces of information in the prompt: micronote, video transcript where the micronote was written, and reference examples from onboarding. The transcript helps to infer abbreviated symbols and fix potential typographical error (D1) (Trajanovski et al., 2021). Reference examples helps to understand users’ writing style (D2).
Note Organization by Theme. To rearrange the note points by emerging themes, we have an atomic order by theme operation (D1, D3). To order by theme, we pass the generated notes to the LLM and instruct it to return the output in a structured, parseable format by using the few-shot prompting technique (Jiang et al., 2023; OpenAI, 2023). We refer users to view the corresponding prompt Figure 11 in the Appendix to learn more about the structure we chose to use.
Figure 4. Note organization (some texts have been cropped for clarity).
Note Modification. To edit individual micronote or the expanded output from LLM and/or the automatically generated theme name (D3), users can click on the [Uncaptioned image] button next to each point.
@@
Summary Generation. To generate a personalized summary, we have a Summarize operation. The generated summary needs to reflect the user’s note as well as the video context (D2). Hence, we passed the full video transcript and expanded note to LLM for the summary.
@@ 5.3. System Walkthrough
We present a hypothetical scenario to illustrate how users might interact with NoTeeline. Kate is preparing a presentation on ‘Ocean Pollution.’ She uses NoTeeline to quickly go through a long 40-minute. She starts by completing the onboarding session and opens a new note named ‘Ocean Pollution Presentation.’ As she starts watching the video, Kate encounters a segment explaining the drastic increase in plastic pollution over the last decade. She quickly types ”plastic pol. $\rightarrow$” as a micronote. Once the video concludes, she presses the [Uncaptioned image] button, and NoTeeline elaborates her note into a detailed point, including the specific statistic: ”Plastic pollution in oceans has increased by 200% in the last 10 years.” This expansion allows Kate to easily integrate precise details into her presentation later.
Now, Kate needs to organize her information into themes so that she can design the slide decks accordingly. She clicks on [Uncaptioned image] and it organizes her notes into three themes - ‘Causes of Ocean Pollution’, ‘Effects on Marine Life’, and ‘Conservation Efforts’. While she is happy with the themes, Kate decides to further personalize the themes. She notices a point ’biodegradable alternatives to plastic’ that would be better suited under ‘Conservation Efforts’. Kate drags it under it and renames the theme to ‘Sustainable Solutions’ to better align with her presentation structure.
Feeling content with the organization of her notes, Kate clicks on [Uncaptioned image] to generate a personalized TL;DR summary for her conclusion slide. Finally, Kate wants to anticipate potential questions from her audience. She opens the [Uncaptioned image] panel. Attempting to answer from memory, Kate correctly addresses the first question but stumbles on the second, prompting her to revisit her notes for a quick recap. Satisfied with her understanding and the comprehensive coverage of her topic, she confidently moves on to finalizing her presentation, equipped with a well-structured and informative slide deck.
@@ 6. Prototype Implementation
NoTeeline is implemented using React.js and is hosted as a web app. All LLM-based operations and text manipulation are performed using OpenAI model, GPT-4-Turbo with hyperparameters set to temperature = 0.5, seed = 1. We chose GPT-4 over other LLMs due to its high-quality output and support for a longer size context window, especially suitable for longer videos (OpenAI, 2023). We use Zustand (Zustand, 2024) to manage the state of NoTeeline by storing user metadata locally.
@@ 7. User Evaluation
We invited 12 participants from our formative interviews to participate in a user study. Participants used both NoTeeline and a baseline to capture notes and evaluate their performance. Please note that participants were not exposed to NoTeeline’s system prior to the user study. They were divided into four subgroups to counterbalance and mitigate any ordering bias. Table 4 shows the distribution of participants across these subgroups. The test session was one hour long.
User Study Procedure. We first gave an overview of study procedures to the participants. Participants were given the NoTeeline web app link to browse from their laptops. They began with the onboarding session where we collected their personalized writing examples (§B.1). After onboarding, participants were given two videos (§B.2) to watch and take notes under the two study conditions. To prevent potential fatigue, participants were given a 5-minute break in between each video. After each video notetaking session, users were asked to complete a questionnaire (Hart and Staveland, 1988; Brooke et al., 1996; Sauro and Dumas, 2009). Finally, we asked participants try out NoTeeline with a video of their choice.
Baseline. To compare NoTeeline with a baseline, i.e., a basic notes app without any automated features, we used a simplified view of our interface by hiding all advanced features (i.e., the synthesis functions described in §5.2.3). This setting mimics the standalone macbook notes app, most of the participants reported to use. We also want to compare the utility between the personalized summary in NoTeeline and the automatically generated summary in existing tools. Hence, we kept the summary feature in the baseline which is similar in nature as NoteGPT (NoteGPT.io, 2024), i.e., a summary that is not conditioned on user notes.
@@ 8. Results and Findings
8.1. NoTeeline helps Users to Write Notes Efficiently with Reduced Disruption
Users took more notes in NoTeeline. The total number of notes is more in NoTeeline ($\mu=18.75,\sigma=5.69$) compared to baseline ($\mu=14.75,\sigma=2.59$) (Figure 6(b)).
Users took a shorter time to write notes in NoTeeline. Compared to baseline ($\mu=27.79,\sigma=8.25$), the notetaking time was 43.9% shorter in NoTeeline ($\mu=15.59,\sigma=5.4$) (Figure 6(c)).
Users faces fewer disruption in watching the video in NoTeeline. Users paused or rewound less with NoTeeline ($\mu=0.56,\sigma=1.86$) compared to baseline ($\mu=4.94,\sigma=5.97$) when they were writing notes (Figure 5).
The results indicate NoTeeline enabled users to capture notes more efficiently by reducing the amount of text they need to write to express their thoughts. This also helps users to deeply concrete without disrupting their focus by pausing or rewinding. P4 directly speaks to this advantage: ‘I felt like I had to.. take every note [in Baseline].. you saw me like pause and rewind… Whereas in [NoTeeline ], I felt like I could just watch it all the way through and not have to rewind.’ The higher frequency of notes also show that NoTeeline does not hinder the notetaking process in any way, rather it helps them to write it with less effort. Similar to how LLM writing tools facilitate more refined thinking (Nakazawa et al., 2022), NoTeeline enables users to engage more actively with video content by allowing them to write notes with less effort.
(a) Average length of note point.
(b) Total count of note point.
(c) Average time per note point.
Figure 6. Comparison of (a) note length, (b) count, and (c) writing time for each participant between NoTeeline and baseline. Majority of the users took higher amount of notes in NoTeeline in less time.
Figure 7. User study results (N=12). Boxplots from left to right: (a) workload measured with NASA TLX (Hart and Staveland, 1988) (7-point Likert scale, lower is better), (b) usability measured with SUS (Brooke et al., 1996) (7-point Likert scale, higher is better), (c) task completion time in seconds (lower is better).
8.2. NoTeeline supports a Wide Range of Notetaking Pattern and Associated Transition
Table 1 presents some example micronotes our participants generated and their corresponding expanded notes generated by NoTeeline.

### v2

@@ 4.3. Design Goals
Based on our needfinding study, iterative prototyping, and evaluation process, we propose three design goals for NoTeeline.
D1. Micronotetaking as the way to take notes from videos to assist rapid writing. The tool should support the quick creation of micronotes and be able to automatically expand it into a full note so that users do not need to write down all the details (C1). Even if the micronote includes grammatical errors or is extremely short, the system should be able to comprehend it. While transforming micronotes, it should be aware of the context, i.e., the topic being discussed in the video. The system should also support iterative editing of micronotes.
D2. Maintain consistency with user’s writing style and preferences to ensure user agency. The system should take into account the specific writing styles of the user and remain consistent with their personal style. At the same time, the final note should exclude information users are not interested in and should not include any hallucinated information (C2).
D3. Each component is linked to a specific stage of the notetaking process to enhance ease of use. The system should link each component to a specific stage in the notetaking process—creation, editing, and revision. It should allow users to engage selectively with these features based on the current stage, providing ease of use throughout the notetaking process (C3)(Pea, 1987; Nielsen, 2006).
5. NoTeeline
Figure 2. The NoTeeline user interface consists of (a) the Cues panel (outlined in blue) for review questions generated based on the notes, (b) the Notes panel (outlined in orange) for user notetaking and note expansion, and (c) the Summary panel (outlined in purple) for note summary. The layout is inspired by the Cornell method, as seen in the side-by-side comparison. The onboarding session, highlighted on the left, collects example notes from the user.
@@ 5.1. Interface Components
We will now describe individual components of NoTeeline and how they facilitate our design goals.
@@
We design an onboarding session to get reference examples of the user specific writing style (D2). Users watch three short clips (20-30 seconds each); and provide a keypoint as well as the corresponding full note for each clip (Figure 2). Once completed, the system will store these examples to use for note expansion afterward. Users need to complete this session only once when they start using NoTeeline for the first time. This session also gives users an initial experience of what a micronote would look like and what kind of full note to expect from the system.
5.1.2. Notetaking Layout
. To better situate the interface around an effective workflow, we took inspiration from the Cornell method to construct the layout (Pauk and Owens, 2010; Saran et al., 2022; Piolat, 2007). We have three main regions – 1) Notes: the region for writing and organizing notes; 2) Cues: auto-generated questions to review the note; and 3) Summary: main takeaway of the note.
We also incorporated the video viewport on the top of the layout so that users can take notes while watching the video without needing to move between different desktop tabs. NoTeeline progressively discloses the features as per users’ current need for note writing, editing, or analyzing (D3). Figure 2 illustrates the notetaking layout and its different components.
Figure 3. Micronote expansion. Users can toggle between the micronote and the expanded note by clicking the Expand/Reduce buttons.
@@ 5.1.2. Notetaking Layout
Micronote Expansion. To construct the full note, we have an atomic expansion operation. For expansion, we utilize these pieces of information in the prompt: micronote, video transcript where the micronote was written, and reference examples from onboarding. The transcript helps to infer abbreviated symbols and fix potential typographical errors (D1) (Trajanovski et al., 2021). Reference examples help to understand users’ writing style (D2). Figure 3 shows an example of micronote expansion.
Note Organization by Theme. To rearrange the note points by emerging themes, we have an atomic order by theme operation. To order by theme, we pass the generated notes to the LLM and instruct it to return the output in a structured, parseable format by using the few-shot prompting technique (Jiang et al., 2023; OpenAI, 2023). We refer users to view the corresponding prompt Figure 14 in the Appendix to learn more about the structure we chose to use. Figure 4 shows an example of how the note organization operation works.
Figure 4. Note organization by themes (some texts have been cropped for clarity). Users can organize the notes by themes or toggle back to the original serial in which they took the notes.
Note Modification. To edit individual micronotes or the expanded output from LLM and/or the automatically generated theme name (D3), users can click on the [Uncaptioned image] button next to each point.
@@
Summary Generation. To generate a personalized summary, we have a Summarize operation. The generated summary needs to reflect the user’s note as well as the video context (D2). Hence, we passed the full video transcript and expanded the note to LLM for the summary.
@@ 5.2. System Walkthrough
We present a hypothetical scenario to illustrate how users might interact with NoTeeline. Kate is preparing a presentation on ‘Ocean Pollution.’ She uses NoTeeline to quickly go through a long 40-minute and scaffold the information for her slide contents. She starts by completing the onboarding session and opens a new note named ‘Ocean Pollution Presentation.’ As she starts watching the video, Kate encounters a segment explaining the drastic increase in plastic pollution over the last decade. She quickly types ”plastic pol. $\rightarrow$” as a micronote. Once the video concludes, she presses the [Uncaptioned image] button, and NoTeeline elaborates her note into a detailed point, including the specific statistic: ”Plastic pollution in oceans has increased by 200% in the last 10 years.” This expansion allows Kate to easily integrate precise details into her presentation later.
Now, Kate needs to organize her information into themes. She clicks on [Uncaptioned image] and it organizes her notes into three themes - ‘Causes of Ocean Pollution’, ‘Effects on Marine Life’, and ‘Conservation Efforts’. While she is happy with the themes, Kate decides to further personalize the themes. She notices a point ’biodegradable alternatives to plastic’ that would be better suited under ‘Conservation Efforts’. Kate drags it under it and renames the theme to ‘Sustainable Solutions’ to better align with her presentation structure.
Feeling content with the organization of her notes, Kate clicks on [Uncaptioned image] to generate a personalized TL; DR summary for her conclusion slide. Finally, Kate wants to anticipate potential questions from her audience. She opens the [Uncaptioned image] panel. Attempting to answer from memory, Kate correctly addresses the first question but stumbles on the second, prompting her to revisit her notes for a quick recap. Satisfied with her understanding and the comprehensive coverage of her topic, she confidently moves on to creating her presentation using her notes from NoTeeline.
@@ 6. Prototype Implementation
NoTeeline is implemented using React.js and is hosted as a web app. All LLM-based operations and text manipulation are performed using the OpenAI model, GPT-4-Turbo with hyperparameters set to temperature = 0.5, seed = 1. We chose GPT-4 over other LLMs due to its high-quality output and support for a longer size context window, especially suitable for longer videos (OpenAI, 2023). We store all the data locally using Chrome localstorage.
@@ 7. User Evaluation
With our user evaluation, we aimed to explore the following research questions:
  1. RQ1.
    Can NoTeeline help users take notes efficiently with reduced disruption?
  2. RQ2.
    Can NoTeeline generate note expansions that align with the user’s micronote and personal writing style?
Participants. We invited 12 participants from our needfinding interviews to participate in a user study. Participants used both NoTeeline and a baseline to capture notes and evaluate their performance. They were divided into four subgroups to counterbalance and mitigate any ordering bias. Table 3 shows the distribution of participants across these subgroups.
User Study Procedure. We first gave an overview of the study procedures to the participants. Participants were given the NoTeeline web app link to browse from their laptops. They began with the onboarding session where we collected their personalized writing examples (§B.1). After onboarding, participants were given two videos (§B.2) to watch and take notes under the two study conditions. To prevent potential fatigue, participants were given a 5-minute break in between each video. After each video notetaking session, users were asked to complete a questionnaire (Hart and Staveland, 1988; Brooke et al., 1996; Sauro and Dumas, 2009). Finally, we asked participants to try out NoTeeline with a video of their choice. Each session was one hour long.
Baseline. To compare NoTeeline with a baseline, i.e., a basic notes app, we used a simplified view of our interface by hiding all advanced features (i.e., the synthesis functions described in §5.1.3). This setting mimics the standalone Macbook Notes app, which most of the participants reported using. We also want to compare the utility between the personalized summary in NoTeeline and the automatically generated summary in existing tools. Hence, we kept the summary feature in the baseline which is similar to NoteGPT (NoteGPT.io, 2024), i.e., a summary that is not conditioned on user notes, rather just summarizes the video transcript.
@@ 8. Results and Findings
8.1. NoTeeline Helps Users Write Notes Efficiently with Reduced Disruption (RQ1)
Users took a shorter time to write notes in NoTeeline. Compared to baseline ($\mu=27.79,\sigma=8.25$), the notetaking time was 43.9% shorter in NoTeeline ($\mu=15.59,\sigma=5.4$) (Figure 6(b)).
Users took more notes in NoTeeline. The total number of notes is more in NoTeeline ($\mu=18.75,\sigma=5.69$) compared to baseline ($\mu=14.75,\sigma=2.59$) (Figure 6(c)).
Users faced fewer disruption in watching the video in NoTeeline. Users paused or rewound less with NoTeeline ($\mu=0.56,\sigma=1.86$) compared to baseline ($\mu=4.94,\sigma=5.97$) when they were writing notes (Figure 5).
The results indicate that NoTeeline enables users to capture notes more efficiently by reducing the amount of text they need to write to express their thoughts. This also helps users to deeply concrete without disrupting their focus by pausing or rewinding. P4 directly speaks to this advantage: ‘I felt like I had to.. take every note [in Baseline].. you saw me like pause and rewind… Whereas in [NoTeeline ], I felt like I could just watch it all the way through and not have to rewind.’ The higher frequency of notes also shows that NoTeeline does not hinder the notetaking process in any way, rather it helps them to write it with less effort.
(a) Average length of note point.
(b) Average time per note point.
(c) Total count of note point.
Figure 6. Comparison of (a) note length, (b) writing time and (c) count for each participant between NoTeeline and baseline. The majority of the users took the higher amount of notes in NoTeeline in less time.
Figure 7. User study results (N=12). Boxplots from left to right: (a) workload measured with NASA TLX (Hart and Staveland, 1988) (7-point Likert scale, lower is better), (b) usability measured with SUS (Brooke et al., 1996) (7-point Likert scale, higher is better)
8.1.1. Workload.
Using the Shapiro-Wilk test of normality, the differences between baseline and NoTeeline for three of the four NASA-TLX questions (Hart and Staveland, 1988) on workload significantly deviate from a Normal distribution. Hence, we compare the differences in workload through a Wilcoxon One-tailed Signed-Rank Test (Neuhäuser, 2011) with a Bonferroni-corrected (Weisstein, 2004) significance level of 0.0125 per condition. Participants reported significantly lower results for Mental Demand, Effort, and Frustration ($p=0.0123,p=0.0024,p=0.0047$). We did not find statistical significance for higher Performance (p = 0.0803). This may be attributed to users’ general familiarity with notetaking as performance was already high in the baseline ($\mu=5.58$ in the baseline versus a slightly higher average of $\mu=6.25$ in NoTeeline on 7-point Likert scale).
8.1.2. Usability.
Usability was measured via an adapted version of SUS (Brooke et al., 1996)) and yielded an overall average of 80.972, which corresponds to an A grade according to the Sauro–Lewis curved grading scale (Sauro and Lewis, 2016). We note that the average baseline SUS values are also high at 80.833. The high baseline usability may be influenced by the improved screen-space optimization in both conditions (C3). At a minimum, our experimental features demonstrate high usability and do not seem to detract from the inherent usability of notetaking methods more familiar to users.
8.2. NoTeeline Generates High Quality Note Expansions (RQ2)
8.2.1. User evaluation.
During the user study, participants evaluate the quality of the automatically generated full note. On average, users rated the expanded note, $\mu=6$ ($\sigma=1.044$) on a scale of 1 to 7 (higher is better) (Figure 10, 1st row).
8.2.2. Linguistic Property.
We report the linguistic quality (in terms of grammaticality, non-redundancy, focus) of notes using GRUEN (Zhu and Bhat, 2020). The result (Table 4 in Appendix) indicates that the points written using NoTeeline ($\mu=0.84,\sigma=0.07$) are of higher quality (i.e: less grammatical error, less redundant and more focused) than the baseline ( $\mu=0.56,\sigma=0.17$) ($p<0.0001,t(11)=10.87$).
8.2.3. Correctness
. To ensure our automatically expanded notes are indeed correctly extracting the information from the transcript and not hallucinating any misinformation (C2), we compute the factual correctness of the generated note using HHEM (Vectara, 2024; Zhang et al., 2023; Azamfirei et al., 2023). We noticed a Hallucination Rate = $6.8\%$ and a Factual Consistency Rate = $93.2\%$. This indicates that our generated notes are correct and consistent with the video content $93.2\%$ of the time.
8.2.4. Consistency with Personal Writing style
To understand if the generated notes in NoTeeline mirror the nuances of a user’s personal writing style, we conduct a stylometric analysis (Stein and zu Eissen, 2007; Li et al., 2023; Argamon et al., 2010) using Mendenhall’s composition characteristic curve (Mendenhall, 1887) and Kilgariff’s Chi-Squared (Kilgarriff, 2001) methods. These methods compare the choice of words and their distribution to infer the writing style. The closer the distribution curves are, the higher their style similarity. In Figure 9, we can see that the word distribution of expanded notes closely aligns with P9’s hand-written note [^2] For space limitation, we randomly show the characteristic curve of one participant., when we utilize the onboarding examples during note generation. The Chi-squared distance (Kilgarriff, 2001) to the handwritten note is $350.90$ and $382.78$, respectively, with and without onboarding (lower is better) (Table 9). The results indicate that NoTeeline achieves a better style match using the provided onboarding examples (Ning et al., 2024).
Participants also self-rated the consistency of expanded notes with their writing styles to be $\mu=5.5$ ( $\sigma=1.19$) on a scale of 1 to 7 (higher is better). P7 mentioned, ‘I have my own writing style… I use symbols like ‘$>$’… [NoTeeline] keeps using my style, so I loved it.
Figure 8. Composition Characteristic curve (Mendenhall, 1887) for P9. The onboarding session examples help NoTeeline generate notes in a similar style that user would write themselves (the lower the gap between two curves, the higher the stylistic similarity)
Expanded
Note
Expanded
Note w/o
onboarding
Relative
Improvement
P1 449.11 423.76 -5.98%
P2 335.21 375.47 +10.72%
P3 342.71 346.82 +1.18%
P4 297.96 336.86 +11.55%
P5 279.94 313.00 +10.56%
P6 423.84 428.98 +1.20%
P7 279.58 309.28 +9.60%
P8 417.74 402.64 -3.75%
P9 376.40 480.91 +21.73%
P10 287.53 353.18 +18.59%
P11 224.38 242.52 +7.48%
P12 496.38 579.93 +14.41%
avg. 350.90 382.78 +8.33%
Figure 9. Chi-squared distance(Kilgarriff, 2001) between the expanded notes from micronote and user’s manual note. Examples of previous note collected during onboarding help us to achieve a better style similarity with the user’s writing. Green and Red denote positive and negative relative performance change, respectively.
8.2.5. Consistency of Micronote and its Corresponding Expanded Note
To ensure our expanded notes are consistent with the main content of the micronotes, we measure the widely used sentence representation similarity, SBERT (Reimers and Gurevych, 2019; Ethayarajh, 2019; Gao et al., 2019). We find that they have a 58% similarity (avg. score = 0.58). We also examined the similarity of the expanded note and the transcript and found it to be 42% (avg. score = 0.42). These results show that the expanded notes are more similar to micronotes than the transcript, indicating that our expanded notes prioritize the micronotes’ contents and do not merely replicate the video transcripts, which will be undesirable.
8.2.6. Consistency of Summary and Note
P1 P2 P3 P4 P5 P6 P7 P8 P9 P10 P11 P12 avg.
Consistency of Micronote, transcript & corresponding expanded note (§8.2.5)
Micronote &
Expanded Note
0.6 0.41 0.73 0.75 0.63 0.58 0.55 0.6 0.48 0.43 0.57 0.63 0.58
Transcript &
Expanded Note
0.47 0.41 0.42 0.42 0.42 0.39 0.42 0.42 0.44 0.46 0.39 0.41 0.42
Consistency of Summary & Note in two conditions (§8.2.6)
NoTeeline 0.87 0.51 0.9 0.63 0.91 0.77 0.67 0.8 0.68 0.7 0.76 0.82 0.75
8.2.7. Error Analysis
Error in LLM generation. GPT-4 sometimes fails to generate the requested note. For example, P7 wrote a micronote to note down her thoughts that are not directly discussed in the video. The model replied with - ‘Please provide the transcript related to the keypoint so I can assist you in creating a note.’. The occurrence of such error is 0.72% (4 out of 559 total calls).
@@ 8.2.7. Error Analysis
8.3. How Users Use NoTeeline and Individual Features
8.3.1. NoTeeline Supports Diverse Notetaking Patterns
Table 2 presents example micronotes our participants generated and their corresponding expanded notes generated by NoTeeline.